# Seat Arrangement Problems under B-utility and W-utility

### v1

@@ 3 Exchange stability (STA)
In this section we study the problem of finding an exchange-stable arrangement, or reporting that none exists. By previous work [4, 6, 9] and Observation 1, we know that, for all three utility types, if agents’ preferences are arbitrary, then determining the existence of an exchange-stable arrangement is NP-complete, even if the seat graph is a matching graph. By contrast, if agents’ preferences are symmetric and the utility type is S-utility, an exchange-stable arrangement always exists, and can be found using the following process: starting from an arbitrary arrangement, while an exchange-blocking pair $\{p,q\}$ exists, switch the seats of $p$ and $q$, and continue until no such pair exists [16, 4]. We now show that this process also works for W-utility.
In order to do that, we need to introduce some new concepts. For an instance of Seat Arrangement with symmetric preferences, let $S$ be the set of different scores of pairs of agents. Given an assignment $\sigma$, $p(\sigma)=\langle t_{1},t_{2},...\rangle$ is a vector of size $|S|$, where $t_{i}$ equals the number of edges in the seat graph $G$ with score equal to the $i$-th highest element in $S$. Let $p(\sigma^{\prime})[i]$ be $t_{i}$, the $i$-th element of $p(\sigma)$. We say that an arrangement $\sigma$ is w-better than an arrangement $\sigma^{\prime}$ if there exists $i$ ($1\leq i\leq|S|$) such that $p(\sigma)[i]<p(\sigma^{\prime})[i]$ and, $\forall j>i$, $p(\sigma)[j]=p(\sigma^{\prime})[j]$.
@@
Let $\sigma$ be an arrangement, with $p$ and $q$ forming an exchange-blocking pair. Let $\sigma(p)=u$ and $\sigma(q)=v$. Furthermore, let $G^{\prime}\subset G$ be the subgraph of $G$ that contains every vertex and edge of $G$ except for $u$, $v$, and their adjacent edges. Let $\sigma_{G^{\prime}}$ be the restriction of $\sigma$ to $G^{\prime}$, so that for agent $r\in P$, if $\sigma(r)\notin\{u,v\}$, then $\sigma(r)=\sigma_{G^{\prime}}(r)$.
@@
Now, let $\sigma^{p\leftrightarrow q}$ be the arrangement in which $p$ and $q$ switch positions with respect to $\sigma$. As the position of every other agent is the same, we have that $\sigma_{G^{\prime}}=\sigma_{G^{\prime}}^{p\leftrightarrow q}$, which means that $p(\sigma_{G^{\prime}})=p(\sigma_{G^{\prime}}^{p\leftrightarrow q})$.
@@
On the other hand, let $G^{\prime\prime}$ be the subgraph of $G$ that includes $u$, $v$, their neighbours, and every edge that includes one of $u$ or $v$. Thus, every edge of $G$ is included in exactly one of $G^{\prime}$ and $G^{\prime\prime}$. As $p$ and $q$ form an exchange blocking pair, their utilities under $\sigma^{p\leftrightarrow q}$ are strictly higher than under $\sigma$. As the utility type is W-utility, this implies that the edge with the lowest score in the neighbourhood of $\sigma_{G^{\prime\prime}}^{p\leftrightarrow q}(p)=v$ has a strictly higher score than the edge with the lowest score in the neighbourhood of $\sigma_{G^{\prime\prime}}(p)=u$, and likewise for $q$. Hence, $p(\sigma_{G^{\prime\prime}}^{p\leftrightarrow q})$ is w-better than $p(\sigma_{G^{\prime\prime}})$. Therefore $p(\sigma^{p\leftrightarrow q})=p(\sigma_{G^{\prime}}^{p\leftrightarrow q})+p(\sigma_{G^{\prime\prime}}^{p\leftrightarrow q})$ is w-better than $p(\sigma)=p(\sigma_{G^{\prime}})+p(\sigma_{G^{\prime\prime}})$, meaning that each swap results in a strictly w-better arrangement. As the number of different arrangements is finite, this process eventually concludes, and the resulting arrangement is exchange-stable. ∎
@@
However, there is no guarantee that this algorithm runs in polynomial time, as in the case of S-utility. In the case of B-utility, if agents’ preferences are symmetric, an exchange-stable arrangement always exists, and can be found in polynomial time. We build an exchange-stable arrangement as follows. First, we order every pair of agents according to its score. Then, we select two agents $p_{i},p_{j}\in P$ such that $\{p_{i},p_{j}\}$ has the highest score, and assign them to neighbouring vertices $u,v$. Every agent that has $p_{i}$ as a favourite agent will be assigned to a neighbour of $u$, and similarly every agent that has $p_{j}$ as their favourite agent will be assigned to a neighbour of $v$, until either all vertices in $N(u)$ (respectively $N(v)$) have an agent assigned to them or all agents that preferred $p_{i}$ (respectively $p_{j}$) the most have been assigned. For all agents $p_{k}$ that have been assigned in the previous step, we allocate to the neighbours of $\sigma(p_{k})$ any agent whose most preferred agent is $p_{k}$. We iterate this process with any new allocation, until we either run out of agents to assign or vertices to be assigned to.
@@
If there are still two neighbouring vertices without agents assigned to them, we repeat the process outlined above, with a new pair $p_{l},p_{m}\in P$ such that $\{p_{l},p_{m}\}$ has the highest score among pairs made out of unassigned agents, and we keep doing this until there is no edge between two unassigned vertices. The remaining agents will be randomly ordered, and assigned to the vertices where they have the highest possible utility according to this order. Algorithm 1 presents this procedure.
@@
$L=\emptyset$triangleright Set of agents allocated to $G$ with a free vertex as a neighbour
@@
create set of ordered pairs $Q$triangleright While there are two vertices sharing an edge without an agent allocated to them
@@
     end $\triangleright$ While there is an unallocated agent whose favourite agent is in $L$
@@
         $\triangleright$ If no free vertex next to $\sigma(s)$, remove $s$ from $L$
@@
Theorem 2.
Proof.
Let $\sigma$ be an arrangement for instance $\sigma$ obtained as the output of Algorithm 1. Let $p,q\in P$ be two different agents such that Algorithm 1 allocated a vertex to $p$ before allocating a vertex to $q$. Assume that $q$ envies $p$. Aiming for a contradiction, assume that $p$ envies $q$ too, which means that they form an exchange-blocking pair. We have that there are three possibilities for when $p$ was allocated a vertex.
The second possibility is that one of the allocated agents $r$ had a vertex in its neighbourhood with no allocated agent, and $r$ was $p$’s most preferred agent in the subset of agents that, at that point in time, were either not allocated or had a free vertex in their neighbourhood. We arrive at a contradiction here because of similar reasons to those of the first point.
We now show the time complexity of Algorithm 1.
@@
Let $I=(P,F,G,B)$ be an instance of Seat Arrangement with symmetric preferences and B-utility. Then Algorithm 1 can be implemented to run in $O(n^{3})$ time, where $n=|P|$, the number of agents in the instance. Under certain conditions, the algorithm can be executed in $O(n^{2})$ time, implying an algorithm whose complexity is linear in the size of the instance.
@@
The third step consists of, for every agent $p$ assigned to a vertex $v$, assigning the agents that like $p$ the most, i.e., the agents in $F_{p}$, to vertices adjacent to $v$, until either all agents in $F_{p}$ are arranged into the graph or all vertices adjacent to $v$ have an agent assigned to them. For every $q\in F_{p}$, either $q$ has already been arranged into the graph, in which case we do nothing, $q$ has not been arranged yet but there is a vertex $u$ adjacent to $v$ with no assigned agent, in which case we arrange $q$ into $u$ (and we repeat this process for all agents in $F_{q}$), or $q$ has not been arranged but there is no such $u$. In this last case, we reduce $c_{q}$ by $1$. If, at some point of the execution, $c_{q}$ reaches 00, it means that agent $q$ cannot be placed next to any of its most-preferred agents, as all of them have been assigned a vertex already and none of these vertices have an adjacent vertex with no agent assigned to it. In this case, we use the pointer from $q$ to the next-best pair $\{q,r\}$ and check all pairs with the same score that also include $q$. For each such $r$, we check if $r$ has been assigned a vertex $w_{1}$, and if there is a free vertex $w_{2}$ adjacent to $w_{1}$. In this case (which corresponds to $r\in L$ in the pseudocode), we assign $q$ to $w_{2}$. If all of the other agents in these pairs have been placed already and $q$ cannot be placed next to them, we move onto the next set of pairs. If these two cases do not hold, then at least some agent $s$ is in one of these pairs with $q$ and has not been assigned a vertex yet (in the pseudocode, such agent $s$ is in $P^{\prime}$). In this case, we count all such agents, store that number in $c_{q}$, create a pointer to the first pair including $q$ after the pairs with the same score as $\{q,s\}$, and add $q$ to every $F_{s}$. Thanks to the use of the different pointers and data structures, for every agent $p$ arranged into the graph, notifying every $F_{p}$ and placing them if possible takes $O(n)$ time. Similarly, for every agent $p$, going down the list of pairs after $c_{p}$ reaches 00 takes $O(n)$ time over the entire execution of the algorithm, because of the pointers linking the pairs that include $p$. Hence, in total, these two steps take $O(n^{2})$ for all agents over the entire execution of the algorithm.
@@
In fact, Algorithm 1 also works for non-symmetric preferences, as long as, for any subset of agents $Q\subseteq P$ , we can always find agents $p_{i},p_{j}\in Q$ such that $p_{i}$ is a favourite agent of $p_{j}$, and vice versa. The only modification that we need to do is that, at the step when the pair with the highest score is chosen, we now choose a pair in which both agents prefer each other the most over all eligible agents.
@@
Let $I=(P,F,G,U)$ be an instance of Seat Arrangement with $1$-dimensional preferences, a path or cycle seat graph, and utility type either S-utility or W-utility. Then the procedure outlines above produces an exchange-stable arrangement in linear time in the number of agents.
@@ 4 Envy-freeness (EFA)
We now present results on determining the existence of envy-free arrangements. We have seen that, generally, determining whether an instance of Seat Arrangement contains an envy-free arrangement is NP-hard, if the utility type is S-utility. An important exception occurs when the seat graph is a matching graph. In that case, EFA is solvable in polynomial time [4].
We now show that EFA with B-utility is NP-hard. In order to prove this result, we are going to reduce from Partition Into Triangles. In this problem, we are given a graph $G=(V,E)$, where $|V|=3n$ for some $n\in\mathbb{Z}$. The goal is to decide whether there exists a partition of $V$ into $3$-element subsets such that the elements of each subset form a triangle in $G$, where a triangle means that each vertex in the subset is a neighbour of every other vertex. This problem is NP-complete, even if $G$ has maximum degree $4$ [5, 17].
@@
Let $I=(P,F,G,B)$ be an instance of Seat Arrangement with a planar seat graph of bounded degree $G$, symmetric and binary agents’ preferences, and B-utility. Then it is NP-complete to decide whether there exists an envy-free arrangement.
@@
We reduce from Partition Into Triangles (PIT). Let $G=(V,E)$ be an instance from PIT with $3n$ vertices, such that every vertex has degree at least $2$ (otherwise no partition exists), and maximum degree $4$. We create an instance $I=(P,F,H,B)$ of the Seat Arrangement problem. For every vertex $v_{i}\in G$ we create an agent $p_{i}\in P$. Let $V_{i}$ be the set of neighbours of $v_{i}$. Then, for $v_{j},v_{k}\in V_{i}$, we create agent $q_{i}^{j,k}\in P$, with $j<k$. This agent represents the possibility of vertex $v_{i}$ being in the same triangle as vertices $v_{j}$ and $v_{k}$. For vertex $v_{i}$, the set of agents of type $q_{i}^{j,k}$ is $Q_{i}\subseteq P$. As $G$ has maximum degree $4$, $v_{i}$ has at most four neighbours in $I$, and as there are six ways of selecting two elements out of a four-element subset, $|Q_{i}|\leq 6$. Finally, for each vertex $v_{i}$ we create several agents, $r_{i}^{l}\in P$, $21\leq l\leq 2\cdot|Q_{i}|+$ 2, which belong to subset $R_{i}\subseteq P$.
For agents $p_{i},q_{i}^{j,k}\in P$, $f_{p_{i}}(q_{i}^{j,k})=f_{q_{i}^{j,k}}(p_{i})=1$. If vertices $v_{i},v_{j},v_{k}$ are all neighbours of each other, $f_{q_{i}^{j,k}}(q_{j}^{i,k})=f_{q_{j}^{i,k}}(q_{i}^{j,k})=1$, and similarly for $q_{i}^{j,k}$ and $q_{k}^{i,j}$, and $q_{j}^{i,k}$ and $q_{k}^{i,j}$. For agents $p_{i},r_{i}^{l}$, $f_{p_{i}}(r_{i}^{l})=f_{r_{i}^{l}}(p_{i})=1$. For every other pair of agents $a,b$, $f_{a}(b)=f_{b}(a)=0$. Finally, the seat graph $H$ consists of $n$ cycles of size $3$, and $3n$ stars, one for each agent. The $i$-th star, corresponding to vertex $i$, has one centre vertex, and $|Q_{i}|+|R_{i}|-1$ leaves. As $1\leq|Q_{i}|\leq 6$ and $4\leq|R_{i}|\leq 14$, every star has between four and nineteen leaves, so $H$ has bounded degree.
Assume there is a partition into triangles. Then if vertices $v_{i},v_{j},v_{k}$ are in one triangle, we assign to one of the cycles of size three agents $q_{i}^{j,k},q_{j}^{i,k},q_{k}^{i,j}$. We do the same for every other triangle. In the $i$-th star, we put agent $p_{i}$ in the middle, and we assign to the leaves the agents from $Q_{i}$ that are not in any cycle, as well as the agents from $R_{i}$. As exactly one agent from $Q_{i}$ is assigned to a vertex in a cycle, and there are $|Q_{i}|+|R_{i}|-1$ leaves, the assignment is valid. It is also an envy-free arrangement, as every agent gets utility $1$, the maximum possible.

### v2

@@ 3 Exchange stability (STA)
In this section we study the problem of finding an exchange-stable arrangement, or reporting that none exists. By previous work [bodlaender_hedonic_2020, cechlarova_complexity_2002, cechlarova_exchange-stable_2005] and Observation 1, we know that, for all three utility types, if agents’ preferences are arbitrary, then determining the existence of an exchange-stable arrangement is NP-complete, even if the seat graph is a matching graph. By contrast, if agents’ preferences are symmetric and the utility type is S-utility, an exchange-stable arrangement always exists, and can be found using the following process: starting from an arbitrary arrangement, while an exchange-blocking pair $\{p,q\}$ exists, switch the seats of $p$ and $q$, and continue until no such pair exists [massand_graphical_2019, bodlaender_hedonic_2020]. In this case, every step of this process produces an arrangement with strictly higher total utility than the previous one, and by a potential function argument it converges on an exchange-stable arrangement. This potential function argument was first introduced by Bogomolnaia and Jackson [bogomolnaia_2002_stability], and has been used since several times [bilo_hedonic_2022, bullinger_2024_topological, massand_graphical_2019, wilczynski_ordinal_2023]. We now show that this process of swapping agents belonging to exchange-blocking pairs also works for W-utility.
In order to do that, we need to introduce some new concepts. For an instance of seat arrangement with symmetric preferences, let $S$ be the set of different scores of pairs of agents. Given an assignment $\sigma$, $n(\sigma)=\langle t_{1},t_{2},$$\rangle$ is a vector of size $|S|$, where $t_{i}$ equals the number of edges in the seat graph $G$ with score equal to the $i$-th highest element in $S$. Let $n(\sigma)[i]$ be $t_{i}$, the $i$-th element of $n(\sigma)$. We say that an arrangement $\sigma$ is w-better than an arrangement $\sigma^{\prime}$ if there exists $i$ ($1\leq i\leq|S|$) such that $n(\sigma)[i]<n(\sigma^{\prime})[i]$ and, $\forall j>i$, $n(\sigma)[j]=n(\sigma^{\prime})[j]$.
@@
Let $\sigma$ be an arrangement, with $q$ and $r$ forming an exchange-blocking pair. Let $\sigma(q)=u$ and $\sigma(r)=v$. Furthermore, let $G^{\prime}\subset G$ be the subgraph of $G$ that contains every vertex and edge of $G$ except for $u$, $v$, and their adjacent edges. Let $\sigma_{G^{\prime}}$ be the restriction of $\sigma$ to $G^{\prime}$, so that for any agent $s\in P$, if $\sigma(s)\notin\{u,v\}$, then $\sigma(s)=\sigma_{G^{\prime}}(s)$.
@@
Now, let $\sigma^{q\leftrightarrow r}$ be the arrangement in which $q$ and $r$ switch positions with respect to $\sigma$. As the position of every other agent is the same, we have that $\sigma_{G^{\prime}}=\sigma_{G^{\prime}}^{q\leftrightarrow r}$, which means that $n(\sigma_{G^{\prime}})=n(\sigma_{G^{\prime}}^{q\leftrightarrow r})$.
@@
On the other hand, let $G^{\prime\prime}$ be the subgraph of $G$ that includes $u$, $v$, their neighbours, and every edge that includes one of $u$ or $v$. Thus, every edge of $G$ is included in exactly one of $G^{\prime}$ and $G^{\prime\prime}$. As $q$ and $r$ form an exchange blocking pair, their utilities under $\sigma^{q\leftrightarrow r}$ are strictly higher than under $\sigma$. As the utility type is W-utility, this implies that the edge with the lowest score in the neighbourhood of $\sigma_{G^{\prime\prime}}^{q\leftrightarrow r}(q)=v$ has a strictly higher score than the edge with the lowest score in the neighbourhood of $\sigma_{G^{\prime\prime}}(q)=u$, and likewise for $r$. Hence, $n(\sigma_{G^{\prime\prime}}^{q\leftrightarrow r})$ is w-better than $n(\sigma_{G^{\prime\prime}})$. Hence $n(\sigma^{q\leftrightarrow r})=n(\sigma_{G^{\prime}}^{q\leftrightarrow r})+n(\sigma_{G^{\prime\prime}}^{q\leftrightarrow r})$ is w-better than $n(\sigma)=n(\sigma_{G^{\prime}})+n(\sigma_{G^{\prime\prime}})$.
Therefore, every time we swap the seats of two agents who form an exchange-blocking pair, we get a strictly w-better arrangement. As the number of different arrangements is finite, this process eventually concludes, and the resulting arrangement is exchange-stable. ∎
However, there is no guarantee that this algorithm runs in polynomial time. In fact, STA with symmetric preferences and W-utility might also be PLS-complete, like STA with symmetric preferences and S-utility. We leave this as an open problem. As an aside, we should also note the similarity between our argument and that of Wilczynski [wilczynski_ordinal_2023], although their way of comparing profiles and our way work differently.
In the case of B-utility, if agents’ preferences are symmetric, an exchange-stable arrangement always exists, by an argument similar to that of Theorem 1. Whilst the algorithm implied by Theorem 1 does not necessarily run in polynomial time, the alternative approach that we now describe will give a polynomial-time algorithm to find an exchange-stable arrangement under B-utility if agents’ preferences are symmetric.
@@
We build an exchange-stable arrangement as follows. First, we order every pair of agents according to its score. Then, we select two agents $p_{i},p_{j}\in P$ such that $\{p_{i},p_{j}\}$ has the highest score, and assign them to neighbouring vertices $u,v$. Every agent that has $p_{i}$ as a favourite agent will be assigned to a neighbour of $u$, and similarly every agent that has $p_{j}$ as their favourite agent will be assigned to a neighbour of $v$, until either all vertices in $N(u)$ (respectively $N(v)$) have an agent assigned to them or all agents that preferred $p_{i}$ (respectively $p_{j}$) the most have been assigned. For all agents $p_{k}$ that have been assigned in the previous step, we allocate to the neighbours of $\sigma(p_{k})$ any agent whose most preferred agent is $p_{k}$. We iterate this process with any new allocation, until we either run out of agents to assign or vertices to be assigned to.
@@
If there are still two neighbouring vertices without agents assigned to them, we repeat the process outlined above, with a new pair $p_{l},p_{m}\in P$ such that $\{p_{l},p_{m}\}$ has the highest score among pairs made out of unassigned agents, and we keep doing this until there is no edge between two unassigned vertices. If there are agents still to be allocated, they can only be assigned to vertices where every neighbour has an agent allocated there. Thus, for every remaining agent $p$ and free vertex $v$ we can calculate the utility that $p$ would get if she is arranged into $v$. Therefore, the last step of the algorithm is to arbitrarily order the remaining agents, and assign each agent to a free vertex where she would get the highest utility among all free vertices. Algorithm 1 presents this procedure.
@@
$L=\emptyset$ $\triangleright$ Set of agents allocated to $G$ with a free vertex as a neighbour
@@
create set of ordered pairs $Q$ $\triangleright$ While there are two vertices sharing an edge without an agent allocated to them
@@
[…]
@@
[…]
@@
Lemma 1 establishes that Algorithm 1 runs in polynomial time, while Theorem 2 establishes that Algorithm 1 produces an exchange-stable matching.
@@
Let $I=(P,F,G,B)$ be an instance of seat arrangement with symmetric preferences and B-utility. Then Algorithm 1 can be implemented to run in $O(n^{3})$ time, where $n=|P|$, the number of agents in the instance. Under certain conditions, the algorithm can be executed in $O(n^{2})$ time, implying that its complexity is linear in the size of the instance.
@@
The third step consists of, for every agent $p$ assigned to a vertex $v$, assigning the agents that like $p$ the most, i.e., the agents in $F_{p}$, to vertices adjacent to $v$, until either all agents in $F_{p}$ are arranged into the graph or all vertices adjacent to $v$ have an agent assigned to them. For every $q\in F_{p}$, either $q$ has already been arranged into the graph, in which case we do nothing, $q$ has not been arranged yet but there is a vertex $u$ adjacent to $v$ with no assigned agent, in which case we arrange $q$ into $u$ (and we repeat this process for all agents in $F_{q}$), or $q$ has not been arranged but there is no such $u$. In this last case, we reduce $c_{q}$ by $1$. If, at some point of the execution, $c_{q}$ reaches 0, it means that agent $q$ cannot be placed next to any of its most-preferred agents, as all of them have been assigned a vertex already and none of these vertices have an adjacent vertex with no agent assigned to it. In this case, we use the pointer from $q$ to the next-best pair $\{q,r\}$ and check all pairs with the same score that also include $q$. For each such $r$, we check if $r$ has been assigned a vertex $w_{1}$, and if there is a free vertex $w_{2}$ adjacent to $w_{1}$. In this case (which corresponds to $r\in L$ in the pseudocode), we assign $q$ to $w_{2}$. If all of the other agents in these pairs have been placed already and $q$ cannot be placed next to them, we move onto the next set of pairs. If these two cases do not hold, then at least some agent $s$ is in one of these pairs with $q$ and has not been assigned a vertex yet (in the pseudocode, such agent $s$ is in $P^{\prime}$). In this case, we count all such agents, store that number in $c_{q}$, create a pointer to the first pair including $q$ after the pairs with the same score as $\{q,s\}$, and add $q$ to every $F_{s}$. Thanks to the use of the different pointers and data structures, for every agent $p$ arranged into the graph, notifying every $F_{p}$ and placing them if possible takes $O(n)$ time. Similarly, for every agent $p$, going down the list of pairs after $c_{p}$ reaches 0 takes $O(n)$ time over the entire execution of the algorithm, because of the pointers linking the pairs that include $p$. Hence, in total, these two steps take $O(n^{2})$ for all agents over the entire execution of the algorithm.
@@
Theorem 2.
Proof.
Let $\sigma$ be an arrangement for instance $I$ obtained as the output of Algorithm 1. Let $p,q\in P$ be two different agents such that Algorithm 1 allocated a vertex to $p$ before allocating a vertex to $q$. Assume that $q$ envies $p$. Aiming for a contradiction, assume that $p$ envies $q$ too, which means that they form an exchange-blocking pair. We have that there are three possibilities for when $p$ was allocated a vertex.
The second possibility is that one of the allocated agents $r$ had a vertex in its neighbourhood with no allocated agent, and $r$ was $p$’s most preferred agent in the subset of agents that, at that point in time, were either not allocated or had a free vertex in their neighbourhood. We arrive at a contradiction here because of similar reasons to those of the first possibility.
The polynomial-time complexity of Algorithm 1 stands in contrast with the PLS-completeness of finding an exchange-stable arrangement under S-utility and symmetric preferences. In fact, similar problems to STA with symmetric preferences and S-utility are also PLS-complete. Some examples are in hedonic games [gairing_2019_computing] and topological distance games [bullinger_2024_topological]. Thus, it is of interest to study whether those other problems become polynomial-time solvable under notions equivalent to B-utility.
@@
Let $I=(P,F,G,U)$ be an instance of seat arrangement with $1$-dimensional preferences, a path or cycle seat graph, and utility type either S-utility or W-utility. Then the procedure outlined above produces an exchange-stable arrangement in linear time in the number of agents.
@@ 4 Envy-freeness (EFA)
We now present results on finding envy-free arrangements or determining that none exists. For all utility types, EFA is in NP, as for any given arrangement we can check for every pair of agents $p,q$ whether $p$ envies $q$. We have seen that, generally, determining whether an instance of seat arrangement contains an envy-free arrangement is NP-complete, if the utility type is S-utility. An important exception occurs when the seat graph is a matching graph. In that case, EFA is solvable in polynomial time [bodlaender_hedonic_2020]. In general, an envy-free arrangement is not guaranteed to exist, so each hardness result in this section implies that the related problem of determining whether an envy-free arrangement exists is NP-complete.
First, we show that EFA with B-utility is NP-hard. In order to prove this result, we are going to reduce from Partition Into Triangles. In this problem, we are given a graph $G=(V,E)$, where $|V|=3n$ for some $n\in\mathbb{Z}$. The goal is to decide whether there exists a partition of $V$ into $3$-element subsets such that the elements of each subset form a triangle in $G$, where a triangle means that each vertex in the subset is a neighbour of every other vertex. This problem is NP-complete, even if $G$ has maximum degree $4$ [caprara2002packing, van_rooij_partition_2013].
@@
Envy-Free Arrangement under B-utility is NP-hard even if agents’ preferences are symmetric and binary, the seat graph has bounded degree, and every graph component is either a cycle or a star.
@@
We reduce from Partition Into Triangles (PIT). Let $G=(V,E)$ be an instance from PIT with $3n$ vertices, such that every vertex has degree at least $2$ (otherwise no partition exists), and maximum degree $4$. We create an instance $I=(P,F,H,B)$ of the seat arrangement problem. For every vertex $v_{i}\in G$ we create an agent $p_{i}\in P$. Let $V_{i}$ be the set of neighbours of $v_{i}$. Then, for $v_{j},v_{k}\in V_{i}$, we create agent $q_{i}^{j,k}\in P$, with $j<k$. This agent represents the possibility of vertex $v_{i}$ being in the same triangle as vertices $v_{j}$ and $v_{k}$. For vertex $v_{i}$, the set of agents of type $q_{i}^{j,k}$ is $Q_{i}\subseteq P$. As $G$ has maximum degree $4$, $v_{i}$ has at most four neighbours in $I$, and as there are six ways of selecting two elements out of a four-element subset, $|Q_{i}|\leq 6$. Finally, for each vertex $v_{i}$ we create several agents, $r_{i}^{l}\in P$, $1\leq l\leq 2\cdot|Q_{i}|+2$, which belong to subset $R_{i}\subseteq P$.
For agents $p_{i},q_{i}^{j,k}\in P$, $f_{p_{i}}(q_{i}^{j,k})=f_{q_{i}^{j,k}}(p_{i})=1$. If vertices $v_{i},v_{j},v_{k}$ are all neighbours of each other, $f_{q_{i}^{j,k}}(q_{j}^{i,k})=f_{q_{j}^{i,k}}(q_{i}^{j,k})=1$, and similarly $f_{q_{i}^{j,k}}(q_{k}^{i,j})=f_{q_{k}^{i,j}}(q_{i}^{j,k})=1$ and $f_{q_{j}^{i,k}}(q_{k}^{i,j})=f_{q_{k}^{i,j}}(q_{j}^{i,k})=1$. For agents $p_{i},r_{i}^{l}$, $f_{p_{i}}(r_{i}^{l})=f_{r_{i}^{l}}(p_{i})=1$. For every other pair of agents $a,b$, $f_{a}(b)=f_{b}(a)=0$. Finally, the seat graph $H$ consists of $n$ cycles of size $3$, and $3n$ stars, one for each agent. The $i$-th star, corresponding to vertex $i$, has one centre vertex, and $|Q_{i}|+|R_{i}|-1$ leaves. As $1\leq|Q_{i}|\leq 6$ and $4\leq|R_{i}|\leq 14$, every star has between four and nineteen leaves, so $H$ has bounded degree.
Assume there is a partition into triangles. Then if vertices $v_{i},v_{j},v_{k}$ are in one triangle, we assign to one of the cycles of size three agents $q_{i}^{j,k},q_{j}^{i,k},q_{k}^{i,j}$. We do the same for every other triangle. In the $i$-th star, we put agent $p_{i}$ in the middle, and we assign to the leaves the agents from $Q_{i}$ that are not in any cycle, as well as the agents from $R_{i}$. As exactly one agent from $Q_{i}$ is assigned to a vertex in a cycle, and there are $|Q_{i}|+|R_{i}|-1$ leaves, the assignment is valid. It is also an envy-free arrangement, as every agent gets utility $1$, the maximum possible. An example of the transformation, with the envy-free arrangement included, is given in Figure 1.
$v_{1}$$v_{2}$$v_{3}$$v_{4}$
$q_{1}^{2,3}$$q_{2}^{1,3}$$q_{3}^{1,2}$$p_{1}$$q_{1}^{2,4}$$q_{1}^{3,4}$$r_{1}^{1}$$r_{1}^{2}$$r_{1}^{3}$$r_{1}^{4}$$r_{1}^{5}$
Figure 1: On the left, subgraph of an example PIT instance, with one triangle from the partition highlighted. On the right, part of an envy-free arrangement, focused on the agents derived from vertex $v_{1}$.